Solve by the elimination method.
0.3x - 0.2y = 4
0.5x + 0.5y = -45/17 (this is a fraction that is -)
I have no idea how to solve this, what do i do with that fraction? Help please.

Answer by stanbon(75887)

Solve by the elimination method.
0.3x - 0.2y = 4
0.5x + 0.5y = -45/17
------------
Multiply thru 1st by 0.5 and 2nd by 0.3 to get:
-------------
0.15x - 0.10y = 2
0.15x + 0.15y =-0.7911
------------------
Subtract 1st from 2nd to solve for "y":
-------
0.05y = -2.7911
y = -5.5882
-------
Substitute into 0.3x-0.2y = 4 to solve for "x":
0.3x -0.2(-5.5882) = 4
0.3x = 2.88235...
x = -9.6078...
Cheers,
Stan H.
==============
Cheers,
Stan H.